Mabel borrowed an amount of money from her father.The loan will be paid back by means of R25000 each every second month for six years,an interest of 7,5% per year,compounded every two months,will be applicable,The amount of loan is


1
Expert's answer
2022-03-29T16:15:16-0400

The present value, PV, of the loan is

"PV=Pmt[\frac{1-\frac{1}{(1+\frac{r}{m})^{nm}}}{\frac{r}{m}}]"

Where Pmt is R25000,

n is 6,

m is 6,

r is 0.075(7.5%)

"\therefore R25000[\frac{1-\frac{1}{(1+\frac{0.075}{6})^{6×6}}}{\frac{0.075}{6}}]"

"PV=R721181.68"

The amount of money Mabel borrowed is R721181.68